OpenStudy (sophadof):

Find the diameter of a sphere with volume 972pi in3.

OpenStudy (anonymous):

vol.=4/3*pi*r^3=972*pi r^3=972*pi/(4/3*pi) r^3=972*3/4=729 r=9